# Examining Inequality in Park Quality for Promoting Health Across 35 Global Cities

### v1

@@ B.1 Determining the Threshold Values for Computing the Linear Models
B.2 Data Cleaning of OSM Tags
@@ B.3 Benchmarking LLM Classifiers
[…]
@@
We tested various settings to see which would deliver the best performance, which was GPT-4, set at a temperature of 0.9, and without providing definitions for tags. To give you a clearer picture, we’ve compiled the results of the top-performing configuration in Table 10. The F1 scores tabulated in the tables show the best results of systematically adjusting the reliability scores for primary and secondary categories from 0 (using any label, regardless of its reliability) to 1 (annotate “none” in all cases). Generally, GPT-4 outperformed its predecessor, GPT-3.5. Adding definitions actually had a negative effect on label quality, possibly due to misleading keywords in the tagging instructions. When it came to the temperature setting, there was no consistent impact, with minimal differences between otherwise equivalent configurations. Interestingly, adding a secondary activity label didn’t improve the quality of the annotation quality (cf. last column of Table 10). In fact, the best results were achieved when the reliability threshold of the secondary annotation was close to 1, rendering all secondary annotations to “none”, thus being equivalent to only using the main activity label. This suggests that the primary labels generated by the system are already of high quality, so putting any weight on a secondary label actually harms the overall score. Based on these findings, we decided to use GPT-4, set at a temperature of 0.9 and without definitions, to label all OSM tags and not impose any threshold on the reliability score.
B.4 Flickr Labels Scoring Model

### v2

@@ B.2 Determining the Threshold Values for Computing the Linear Models
B.3 Data Cleaning of OSM Tags
@@ B.4 Benchmarking LLM Classifiers
$TP^{\prime}={TP^{i}_{main}\cdot reliability^{i}+TP^{i}_{2nd}\cdot(1-reliability^{i})}\hfill\text{(true positives)}$ (4)
$FP^{\prime}={FP^{i}_{main}\cdot reliability^{i}+FP^{i}_{2nd}\cdot(1-reliability^{i})}\hfill\text{(false positives)}$ (5)
$FN^{\prime}={FN^{i}_{main}\cdot reliability^{i}+FN^{i}_{2nd}\cdot(1-reliability^{i})}\hfill\text{(false negatives)}$ (6)
$reliability^{i}=\frac{mean(reliability^{i}_{main})}{mean(reliability^{i}_{main})+mean(reliability^{i}_{2nd})}\hfill\text{(Ratio of reliability between main and secondary category)}$ (7)
@@
We tested various settings to see which would deliver the best performance, which was GPT-4, set at a temperature of 0.9, and without providing definitions for tags. To give you a clearer picture, we’ve compiled the results of the top-performing configuration in Table 8. The F1 scores tabulated in the tables show the best results of systematically adjusting the reliability scores for primary and secondary categories from 0 (using any label, regardless of its reliability) to 1 (annotate “none” in all cases). Generally, GPT-4 outperformed its predecessor, GPT-3.5. Adding definitions actually had a negative effect on label quality, possibly due to misleading keywords in the tagging instructions. When it came to the temperature setting, there was no consistent impact, with minimal differences between otherwise equivalent configurations. Interestingly, adding a secondary activity label didn’t improve the annotation quality (cf. last column of Table 8). In fact, the best results were achieved when the reliability threshold of the secondary annotation was close to 1, rendering all secondary annotations to “none”, thus being equivalent to only using the main activity label. This suggests that the primary labels generated by the system are already of high quality, so putting any weight on a secondary label actually harms the overall score. Based on these findings, we decided to use GPT-4, set at a temperature of 0.9 and without definitions, to label all OSM tags and not impose any threshold on the reliability score.
B.5 Modeling Average Park Offerings
To model average park offerings, we used independent regression models for park elements and spaces. During the computation of the regression models, parks with very low activity counts in a specific category were excluded. This exclusion was necessary to prevent artificially flattening the regression lines due to close-to-zero values, which would distort the normalization. The specific threshold for excluding parks with low activity counts was determined empirically by analyzing the histograms of the values. This approach enabled us to identify an appropriate cutoff point for excluding parks with insufficient activity data, ensuring the reliability of the regression models. For a visual representation of the exclusion process and the determination of the threshold, refer to Figure 9 and Figure 10.
B.6 Orthogonality of Park Elements and Spaces
Park elements include points of interest, individual trees, benches, and similar items. Park spaces, however, include areas like forests, sports fields, and buildings. There can be cases where a park area is broken down into its individual parts, like a playground with separately mapped features like swings, slides, or spinning equipment. But these cases are pretty rare in OSM mapping. Likewise, unless a tree is particularly important, areas tagged as natural=wood should not include individual trees according to the mapping guidelines[185]. Based on these observations, we hypothesized that it would be acceptable to combine scores from park features and areas linearly, as they contribute differently to the overall offering of facilities for health-promoting activities.
Activity Category Mean Pearson Correlation
Physical 0.16
Nature-appreciation 0.05
Environmental 0.18
Social 0.06
Cultural 0.18
B.7 Semantic Matching of Flickr Labels and OSM Tags
In our validation, we used a global dataset of geotagged photos from Flickr, from which we selected all 10,711,513 images that were taken within one of the parks from 35 cities in our study. These images came with user-generated labels and were also partially annotated with computer vision labels from a computer vision algorithm[57, 58]. To obtain semantically equivalent representations of Flickr labels and OSM tags, we employed Sentence-BERT (S-BERT[186]) for text embeddings. We formulated this task as an asymmetric semantic search problem, where the Flickr label was the search term, and the goal was to find the closest matching OSM tag. Given the worldwide reach of our study, the multiple languages present in the user-generated Flickr labels created a challenge in mapping them to the corresponding OSM tags, which were all in English. To address this, we identified the top three languages besides English used in the tags of each city, using the Google MediaPipe[187] Language Detection Model[188]. To ensure that the language detection was accurate and to eliminate named entities, we only used labels where the language detection indicated a confidence of 50% or more. We then translated those tags to English using the respective OPUS machine translation models[189].
A detailed review of the matched pairs revealed that, as anticipated, most pairings were logical based on the text similarity between labels and tags with definitions. However, some minor adjustments were still needed, as some matches were not entirely consistent with the theme of health-promoting activities in parks. For example, the term “outdoor” was initially linked to swimming_pool=outdoor. But as there cannot be a suitable equivalent for “outdoor” on OSM, we removed this pairing and equivalent ones, such as “park,” as all photos were taken in parks. Another instance was the pairing of “water”, which did not capture the specific role of water features in parks in promoting health. We manually adjusted this to water=river, which better reflects bodies of water commonly found in parks. Through this review step, we improved the quality of the matched pairs, ensuring they more closely align with the theme of health-promoting park activities. The need for this manual step should not diminish the effectiveness of the semantic search within sentence embeddings. It was merely to eliminate labels that could not meaningfully correspond with an OSM tag and to match a few labels with more domain-relevant tags. This matching process yielded 2,171 label-tag pairs in total. Of these, 1,432 pairs corresponded to an OSM tag with health-promoting benefits, such as “steeplechase” being matched to athletics=steeplechase involving physical health benefits, while 739 pairs, such as “Lamp Post” being matched to man_made=lamp_post did not imply health benefits.
Having assured that the matchings are accurate, we proceeded to profile the parks based on the activities associated with the matched OSM tags, following the same scoring approach as what we used for the OSM park elements and spaces (cf. Equation 2). In our validation, we chose a minimum of 250 images from each park and at least 15 parks in each city. This criterion was established to secure a robust number of images for each park, enhancing the accuracy of our analysis. This approach helped us avoid any potential bias that could have been introduced by individual photographers if a park had only a few images.
@@ B.7 Semantic Matching of Flickr Labels and OSM Tags
B.8 User Studies with Urban Designers and Park Maintenance Experts
We conducted three semi-structured interviews, each lasting about 30 minutes, with urban designers and park maintenance experts. Participants were recruited through our network and by direct email to municipalities. The experts’ backgrounds were as follows:
  • Lecturer and urban designer (E1). 20 years of practical and academic experience in Hong Kong and London.
  • Urban designer and master planner (E2). 30 years of experience in the United States, United Kingdom, Europe, and the Middle East.
  • Parks development manager at a local council (E3). 12 years of experience as conversationalist, arboricultural manager in diverse contexts.
The online interviews consisted of 4 steps.
  1. 1.
    Introduction. Three questions to understand their experience in urban park design and management.
  2. 2.
    Study brief. A summary of our study and main findings.
  3. 3.
    Visualization. Interacting with the visualization and exploring a chosen city.
  4. 4.
    Debriefing. Open-ended questions and feedback on the project.
We qualitatively analyzed the interview transcripts [190, 191] (recorded with consent) to understand the implications of our work for professional practice, verify inter-cultural validity, and to further validate the taxonomy of activities.